# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v11

@@ 4.2 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
$It=\left\langle ot,{Tt}_{h},S_{ot},M_{I},t_{1},S_{M_{I}}\right\rangle$ (4.7)
@@
Here, the ontology $ot$ refers to the object containing knowledge or samples; ${Tt}_{h}$ is the time at which its state is formed; and $S_{ot}(ot,{Tt}_{h})$ is a set of business-oriented semantic assertions or facts about the ontology. $M_{I}$ functions similarly to the hardware interface and API software of $M$, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is the physical realization of $S_{ot}(ot,{Tt}_{h})$ via an enabling mapping, ensuring that the information can be machine-read and entered into $M$.
@@ 4.3 Formal Modeling and Implementation of Model Learning
Model learning for $M$ involves fusing its original information $Io$ with the training information $It$, resulting in updated information:
$Iu=\left\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\right\rangle$ (4.10)
@@
indicates that the input state $S_{M_{I}}\left(M_{I},t_{1}\right)$, which was learned by $M$ at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of $M$ at $t_{2}$.
$\begin{split}{\varphi u}_{2}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))))\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.13)
shows that after the first learning step, $M$ remains capable of further learning and optimization.
${\varphi u}_{3}=$ $State^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\wedge Input^{3}\left(M,t_{2},S_{M_{I}}(M_{I},t_{2})\right)$ (4.14)
$\wedge Processable^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)\rightarrow\exists f_{process}^{2}\left(M,S_{M_{I}}(M_{I},t_{2})\right)$
$\left(Eq^{2}\left(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))\right)\wedge Output^{3}\left(M,t_{3},S_{M_{O}}(M_{O},t_{3})\right)\right)$
indicates that when the training information $It$ is processable by $M$, it can also produce output information.
This state is then enabled and mapped to the physical state of the carrier $M$ at time $t_{2}$:
$S_{M}\left(M,t_{2}\right)=\left\{{\psi u}_{1},{\psi u}_{2},{\psi u}_{3}\right\}$ (4.15)
@@ 4.4 User Input Information for the Model
User input information is formally given by:
$Iq=\left\langle oq,{Tq}_{h},S_{oq},M_{I},t_{2},S_{M_{I}}\right\rangle$ (4.16)
@@
Here, the ontology $oq$ represents the user’s intention or device, with ${Tq}_{h}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,{Tq}_{h})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,{Tq}_{h})$ onto the model’s interface device $M_{I}$.
@@ 4.5 Information Processing, Output, and Online Learning of the Model
When the model $M$ processes input and produces output, the information it carries can be represented as:
$Iv=\left\langle M\cup M_{I}\cup M_{O},t_{2}\cup t_{3},S_{M\cup M_{I}\cup M_{O}},M,t_{3},S_{M}\right\rangle$ (4.19)
@@
$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2},{\varphi v}_{3}\right\}$ (4.20)
${\varphi v}_{1}={State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)$ (4.21)
indicates that $\Phi^{2}(M,t_{2})$ has been assigned to $M$ at time $t_{2}$.
$\begin{split}{\varphi v}_{2}=&{Input}^{3}(M,t_{2},S_{M_{I}}\left(M_{I},t_{2}\right))\land{Eq}^{2}(S_{M_{O}}\left(M_{O},t_{3}\right),f_{process}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right)))\\ &\land{Output}^{3}(M,t_{3},S_{M_{O}}\left(M_{O},t_{3}\right))\end{split}$ (4.22)
indicates that when the input state $S_{M_{I}}\left(M_{I},t_{2}\right)$ is received at $t_{2}$, $M$ produces the output state $S_{M_{O}}\left(M_{O},t_{3}\right)$ at $t_{3}$.
$\begin{split}{\varphi v}_{3}=&{State}^{3}\left(M,t_{2},\Phi^{2}(M,t_{2})\right)\land{Input}^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land{Learnable}^{2}(M,S_{M_{I}}\left(M_{I},t_{2}\right))\\ &\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))({Eq}^{2}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))))\\ &\land{State}^{3}\left(M,t_{3},\Phi^{2}(M,t_{3}))\right)\end{split}$ (4.23)
expresses that if the input information can be learned by $M$, then $M$ can achieve online learning while processing and outputting information.
The physical realization of this state through an enabling mapping is:
$S_{M}(M,t_{3})=\left\{{\psi v}_{1},{\psi v}_{2},{\psi v}_{3}\right\}$ (4.24)
@@
$S_{M\cup M_{I}}\left(M\cup M_{I},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.26)
@@
The formal expression of the physical state of the model’s output device $M_{O}$ at time $t_{3}$ is:
$S_{M_{O}}\left(M_{O},t_{3}\right)=\left\{{\psi r}_{1}\right\}$ (4.28)
The model can repeat the processes described in Sections 3.2 to 3.5, iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
4.6 Multi-Stage Information Causal Chains
From systems engineering perspective, the realization of information in machine learning involves more than just five information mappings. In practice, stages such as requirements analysis, design and development, and system integration are also necessary. Specifically, in the four stages of model construction, learning and training, problem input, and answer output, and according to the corollary on information chains in Section 3.3, there is a causal and progressive process that moves from conceptual ideas to mind maps, program code, and finally to physical implementation. In this process, the earlier content serves as the ontological state, while the subsequent content becomes the carrier state, sequentially forming conceptual information, design information, and implementation information (see Fig.2).
Fig.2 presents a comprehensive framework of formal expressions and causal information progression for each stage of the machine learning process, from the perspectives of OIT and finite automaton state implementation. This framework transcends specific algorithms or models, offering a novel, unified, and general framework—MLT-MF—for addressing several key issues in the field of machine learning. In Appendix C, we further provide an example of its application to feedforward neural networks, thereby testing the correctness and applicability of the framework and supporting the verification of related theorems.
Figure 2: Causal Chains of Multiple Information Mappings in Machine Learning. In machine learning, the four distinct stages—model construction, learning and training, problem input, and answer output—each encompass four levels: conceptual ideas, mind maps, program code, and physical models. Generally, the content within the same level shares a common semantic logic and can thus be represented using a unified formal system. Furthermore, owing to the transmissibility of information, the innermost physical model can be regarded as the result of an enabling mapping from the outer conceptual ideas or mind maps. The output stage can, in turn, be viewed as a reverse causal transmission from the physical model back to conceptual ideas. This demonstrates that the system comprises multiple stages, with each stage containing various combinatorial information causal chains.
@@ 5.1 Theorem of Model Interpretability
According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model (Glanois et al., 2024), the following concepts are established based on the MLT-MF framework:
@@
(1) For any $1\leq i\leq n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$ and $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$ must hold.
@@
Proof: For any $1\leq i\leq n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let $S_{on}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{op}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 4, there exist information objects $Ig(t_{i})=\langle oq(t_{i}),Tq_{h}(t_{i}),S_{op},M_{I},t_{i},S_{M_{I}}\rangle$ and $Ir(t_{i+1})=\langle or(t_{i+1}),Tr_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\rangle$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{on}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i})=\left\langle ou(t_{i}),{Tu}_{h}(t_{i}),S_{ou},M,t_{i},S_{M}\right\rangle$, $Iq(t_{i})=\left\langle oq(t_{i}),{Tq}_{h}(t_{i}),S_{oq},M_{I},t_{i},S_{M_{I}}\right\rangle$, and $Ir(t_{i+1})=\left\langle or(t_{i+1}),{Tr}_{h}(t_{i+1}),S_{or},M_{O},t_{i+1},S_{M_{O}}\right\rangle$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\quad\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$. Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.

### v12

@@ 4.2 Formal Modeling and Realization of Training Information
The second stage in Fig. 2 is the realization of training information. According to the general formalization of information, its formal expression is straightforward:
$It:S_{ot}(ot,Tt_{h})\rightrightarrows S_{M_{I}}(M_{I},t_{1})$ (4.7)
@@
Here, the ontology $ot$ refers to the object containing knowledge or samples; ${Tt}_{h}$ is the time at which its state is formed; and $S_{ot}(ot,{Tt}_{h})$ is a set of business-oriented semantic assertions or facts about the ontology. $M_{I}$ functions similarly to the hardware interface and API software of $M$, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}\left(M_{I},t_{1}\right)$ is the physical realization of $S_{ot}(ot,{Tt}_{h})$ via an enabling mapping, ensuring that the information can be machine-read and entered into $M$. In practice, much training information falls under the noisy information defined in Definition 3. The structure of noisy information will affect the effectiveness of model training.
@@ 4.3 Formal Modeling and Implementation of Model Learning
In the third stage of Fig. 2, $M$’s learning is precisely the fusion of the original information $Io$ with the training information $It$, optimizing it into updated information:
$Iu:S_{M\cup M_{I}}(M\cup M_{I},t_{1})\rightrightarrows S_{M}(M,t_{2})$ (4.10)
@@
indicates that the input state $S_{M_{I}}\left(M_{I},t_{1}\right)$, which was learned by $M$ at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of $M$ at $t_{2}$. As shown in Equations (4.3) and (4.4), $\varphi_{u_{2}}$ and $\varphi_{u_{3}}$ are logically identical to $\varphi_{o_{2}}$ and $\varphi_{o_{3}}$, respectively, differing only in the underlying ontological object and time. This indicates that the state retains the capacity for continued learning and processing even after the initial learning and optimization phases.
@@ 4.4 User Input Information for the Model
User input information in the fourth stage in Fig. 2 is formally given by:
$Iq:S_{oq}(oq,Tq_{h})\rightrightarrows S_{M_{I}}(M_{I},t_{2})$ (4.13)
@@
Here, the ontology $oq$ represents the user’s intention or device, with ${Tq}_{h}$ denoting the time at which the ontological state is formed. The state $S_{oq}(oq,{Tq}_{h})$ is the formal representation of the user’s needs. The carrier state $S_{M_{I}}(M_{I},t_{2})$ is the physical realization, via enabling mapping, of $S_{oq}(oq,{Tq}_{h})$ onto the model’s interface device $M_{I}$. User input information is also typically affected by noise. Its noisy information structure will be directly related to the $M$’s output error.
@@ 4.5 Information Processing, Output, and Online Learning of the Model
In the fifth stage shown in Fig. 2, when the model $M$ processes input and produces output, the information it carries can be represented as:
$Iv:S_{M\cup M_{I}\cup M_{O}}(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3})\rightrightarrows S_{M}(M,t_{3})$ (4.16)
@@
$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\cup t_{3}\right)=\left\{{\varphi v}_{1},{\varphi v}_{2}\right\}$ (4.17)
$M_{O}$ is $M$’s output device. The formulas $\varphi_{v_{1}}$ and $\varphi_{v_{2}}$ are logically identical to $\varphi_{u_{2}}$ and $\varphi_{u_{3}}$, respectively, differing only in their temporal references. This demonstrates that if the input information is processable and learnable by $M$, then $M$ can perform online learning concurrently with information processing and output operations.
@@
$S_{M\cup M_{I}\cup M_{O}}\left(M\cup M_{I}\cup M_{O},t_{2}\right)=\left\{{\varphi r}_{1}\right\}$ (4.19)
@@
Finally, it should be emphasized that $M$ can repeat the processes described in Sections 4.2 to 4.5, iterating through cycles of training and information processing to meet user demands and continuously improve its performance.
@@ 5.1 Theorem of Model Interpretability
According to the view that interpretability should be a concept dependent on the domain of discourse and the corresponding logical system, and that interpretability should also be related to the specific inputs and outputs of the model [16], the following concepts are established based on the MLT-MF framework:
@@
(1) For any $i=1,\dots,n-1$, a necessary and sufficient condition for the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ to be interpretable in the formal system $\mathcal{L}$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is that there exist $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$ and $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ representing the information carried by $M$ and its input device $M_{I}$ at time $t_{i}$, respectively, and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$ representing the information carried by the output device $M_{O}$ at time $t_{i+1}$. The ontological states of these three pieces of information must each correspond to an interpretation in $\mathcal{L}$, and $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$ must hold.
@@
Proof: For any $i=1,\dots,n-1$, suppose that the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. According to Definition 4, there exist three interpretations $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$ in the formal system $\mathcal{L}$, respectively over the finite domains $D_{M}$, $D_{M_{I}}$, and $D_{M_{O}}$, such that $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$, and $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are each the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively.
Let $S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))=E_{M}$, $S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))=E_{M_{I}}$ and $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))=E_{M_{O}}$. Then, according to Definition 4, there exist information objects $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$, $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$ such that $S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),Tq_{h}(t_{i})))$. Thus, the necessity of the first part of the theorem is established.
On the other hand, for any $1\leq i\leq n-1$, suppose there exist information objects $Iu(t_{i}):S_{ou}(ou(t_{i}),Tu_{h}(t_{i}))\rightrightarrows S_{M}(M,t_{i})$, $Iq(t_{i}):S_{oq}(oq(t_{i}),Tq_{h}(t_{i}))\rightrightarrows S_{M_{I}}(M_{I},t_{i})$ and $Ir(t_{i+1}):S_{or}(or(t_{i+1}),Tr_{h}(t_{i+1}))\rightrightarrows S_{M_{O}}(M_{O},t_{i+1})$, such that in the formal system $\mathcal{L}$, $S_{or}(or(t_{i+1}),{Tr}_{h}(t_{i+1}))\subseteq Cn(S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))\cup S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i})))$. Then, according to Lemma A.1, $E_{M}=S_{ou}(ou(t_{i}),{Tu}_{h}(t_{i}))$, $E_{M_{I}}=S_{oq}(oq(t_{i}),{Tq}_{h}(t_{i}))$, and $E_{M_{O}}=S_{or}\left(or\left(t_{i+1}\right),{Tr}_{h}\left(t_{i+1}\right)\right)$ are interpretations of $\mathcal{L}$ over the three finite domains $ou(t_{i})\times{Tu}_{h}(t_{i})$, $oq(t_{i})\times{Tq}_{h}(t_{i})$, and $r(t_{i+1})\times{Tr}_{h}(t_{i+1})$, respectively, and $E_{M_{O}}\subseteq Cn(E_{M}\cup E_{M_{I}})$. At the same time, $S_{M}\left(M,t_{i}\right)$, $S_{M_{I}}\left(M_{I},t_{i}\right)$, and $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ are the enabling mappings of $E_{M}$, $E_{M_{I}}$, and $E_{M_{O}}$, respectively. By Definition 4, the output $S_{M_{O}}\left(M_{O},t_{i+1}\right)$ of model $M$ under the input $S_{M_{I}}\left(M_{I},t_{i}\right)$ is interpretable in the formal system $\mathcal{L}$. Therefore, the sufficiency of the first part of the theorem is established.